# Ratio asymptotics and zero density for orthogonal polynomials with varying Verblunsky coefficients

### v1

@@ 2.2. CMV matrices
From the spectral theory point of view, the probability measure $\mu$ associated with the sequence of Verblunsky coefficients $\{\alpha_{n}\}_{n=0}^{\infty}\subset{\mathbb{D}}^{\infty}$ is the spectral measure with respect to $e_{0}$ of the unitary operator on $\ell^{2}(\mathbb{N}_{0})$, referred to as the CMV matrix:
$\mathcal{C}=\begin{pmatrix}\bar{\alpha}_{0}&\bar{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&0&0&\dots\\ \rho_{0}&-\bar{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&0&0&\dots\\ 0&\bar{\alpha}_{2}\rho_{1}&-\bar{\alpha}_{2}\alpha_{1}&\bar{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&\dots\\ 0&\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\bar{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&\dots\\ 0&0&0&\bar{\alpha}_{4}\rho_{3}&-\bar{\alpha}_{4}\alpha_{3}&\dots\\ \dots&\dots&\dots&\dots&\dots&\dots\end{pmatrix}$, (2.6)
@@
Given a CMV matrix $\mathcal{C}$ associated with Verblunsky coefficients $\{\alpha_{n}\}_{n=0}^{\infty}$, the once-stripped CMV matrix, denoted $\mathcal{C}^{(1)}$, is the CMV matrix constructed from the shifted sequence of Verblunsky coefficients $\{\alpha_{n}\}_{n=1}^{\infty}$; that is, the sequence obtained by removing $\alpha_{0}$ and reindexing. Similar definition applies to $\mathcal{C}^{(j)}$, which is the CMV matrix constructed from $\{\alpha_{n}\}_{n=j}^{\infty}$, for any $j\in\mathbb{N}_{0}$.
Let us also introduce the cut-off CMV matrix $\mathcal{C}_{n}$, which is the top-left $n\times n$ submatrix of $\mathcal{C}$. It is easy to show that $\Phi_{n}$ is the characteristic polynomial of $\mathcal{C}_{n}$, i.e,
@@
Inspecting the structure of the CMV matrix shows that $\mathcal{C}_{n}$ depends only on the first $n$ Verblunsky coefficients. When needed, we will indicate this dependence by writing
@@
The paraorthogonal polynomials (2.5) therefore satisfy
@@ 2.3. Carathéodory, Schur, and $m$-functions
$F(z)=\int_{0}^{2\pi}\frac{e^{{\rm i}\theta}+z}{e^{{\rm i}\theta}-z}{\rm d}{\mu(e^{{\rm i}\theta})}=\langle e_{0},(\mathcal{C}+z)(\mathcal{C}-z)^{-1}e_{0}\rangle,\quad z\in{\mathbb{D}}.$ (2.10)
@@
The Schur function is analytic $\mathbb{D}$} and satisfies $\sup_{z\in{\mathbb{D}}}|f(z)|\leq 1$. It will be convenient for us to extend (2.12) and also (2.16) below to $z\in{\mathbb{C}}\setminus\mathop{\mathrm{supp}}\nolimits\,\mu$, which requires allowing $f$ to take value $\infty$ at those points $z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}$ where F(z)=1F(z)=-1. For instance, under this convention, for the free CMV matrix ($\alpha_{n}\equiv 0$) we obtain $f(z)=0$ for all $z\in{\mathbb{D}}$ and $f(z)=\infty$ for all $z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}$, see Section 2.4. From (2.11), we obtain
@@
Finally, it will be convenient to introduce notation for the $(0,0)$-entry of the resolvent of $\mathcal{C}$:
$m(z)=\int_{0}^{2\pi}\frac{{{\rm d}{\mu(e^{{\rm i}\theta})}}}{z-e^{{\rm i}\theta}}=\langle e_{0},(z-\mathcal{C})^{-1}e_{0}\rangle,\quad z\in{\mathbb{C}}\setminus\mathop{\mathrm{supp}}\nolimits\,\mu.$ (2.14)
@@
Formulas (2.15)–(2.17) are immediate from the definitions. Next, since $f$ does not take infinite values in $\overline{{\mathbb{D}}}\setminus\mathop{\mathrm{supp}}\nolimits\mu$, it follows from (2.13) that $f(z)\neq 0$ for $z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}$ as well as for $z\in{\partial\mathbb{D}}\setminus\mathop{\mathrm{supp}}\nolimits\mu$. Combining this with the first equality in (2.16) produces the claim (ii). ∎
@@ 2.4. Constant Verblunsky coefficients
In the so-called free case when $\alpha=0$, we have $\Phi_{n}(z)=z^{n}$ and the corresponding measure of orthogonality is the Lebesgue measure on $\partial\mathbb{D}$}. For $\alpha\in{\mathbb{D}}\setminus\{0\}$, the orthogonal polynomials $\Phi_{n}$ are known as the Geronimus polynomials, see [geronimus_vkgu66], as well as [golinski_maa99, Sec. 2] or [simon_opuc1, Ex. 1.6.12].
@@
where $\mathop{\mathrm{diag}}\nolimits$ denotes a diagonal matrix. The spectral measure $\mu$ of $\mathcal{C}(\alpha)$ with respect to $e_{0}$ is the Dirac delta measure $\delta_{\bar{\alpha}}$ supported on the one-point set $\{\bar{\alpha}\}$. This measure is trivial, and therefore it does not uniquely determine its orthogonal polynomials. Nevertheless, all the functions discussed below are unambiguously defined and consistent with this setting $\alpha\in{\partial\mathbb{D}}$, $\mu=\delta_{\bar{\alpha}}$, and
$\Phi_{n}{(z)=(z-\bar{\alpha})(z+1)^{n-1}},$ (2.18)
$\Phi_{n}^{(\beta)}{(z)=(z-\bar{\alpha})(z+1)^{n-2}(z+\bar{\beta}\alpha)},$ (2.19)
see (2.7) and (2.8).
We denote $a=|\alpha|\in[0,1]$,
@@
$\Gamma_{a}=\left\{e^{{\rm i}\theta}\in{\partial\mathbb{D}}\;\big|\;\theta_{a}\leq\theta\leq 2\pi-\theta_{a}\right\}.$ (2.21)
@@
$\sigma_{\alpha}=\begin{cases}\Gamma_{a}&\mbox{ if }|\alpha+\tfrac{1}{2}|\leq\tfrac{1}{2},\\[2.0pt] \Gamma_{a}\cup\{\frac{1+\bar{\alpha}}{1+\alpha}\}&\mbox{ if }|\alpha+\tfrac{1}{2}|>\tfrac{1}{2}.\end{cases}$ (2.22)
@@
$\sqrt{\left(z-e^{{\rm i}\theta_{a}}\right)\left(z-e^{-{\rm i}\theta_{a}}\right)}=\sqrt{(z-1)^{2}+4za^{2}},$ (2.23)
@@
$\lim_{z\to 0}\sqrt{\left(z-e^{{\rm i}\theta_{a}}\right)\left(z-e^{-{\rm i}\theta_{a}}\right)}=1.$ (2.24)
@@
$G_{a}(z)=\frac{1}{2}\left(z+1+\sqrt{\left(z-e^{{\rm i}\theta_{a}}\right)\left(z-e^{-{\rm i}\theta_{a}}\right)}\right).$ (2.25)
@@
    $G_{0}(z)=\begin{cases}1,&\mbox{ if }z\in{\mathbb{D}},\\ z,&\mbox{ if }z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}},\end{cases}\qquad$ $F_{0}(z)=\begin{cases}1,&\mbox{ if }z\in{\mathbb{D}},\\ -1,&\mbox{ if }z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}},\end{cases}$ (2.28)
    $m_{0}(z)=\begin{cases}0,&\mbox{ if }z\in{\mathbb{D}},\\ \frac{1}{z},&\mbox{ if }z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}},\end{cases}\qquad$ $f_{0}(z)=\begin{cases}0,&\mbox{ if }z\in{\mathbb{D}},\\ \infty,&\mbox{ if }z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}.\end{cases}$ (2.29)
@@
Part (iii): The expression for $F_{\alpha}$ can be found in [golinski_maa99, Eq. (2.5)] (note that our $\alpha$ and $G_{a}$ correspond to their $\bar{a}$ and $z_{1}$, respectively). Combining this formula for $F_{\alpha}$ with (2.14) implies (2.31). Finally, expression (2.32) appears in [simon_opuc1, Eq. (1.6.82)]. Alternatively, one can deduce (2.32) by employing (2.12) and (2.27). ∎
@@ 2.5. Periodic OPUC
$\mathcal{B}_{\Delta}=\Delta^{-1}([-2,2])$ (2.35)
@@
In particular, the case of constant Verblunsky coefficients $\alpha_{n}\equiv\alpha$ from the previous section corresponds to
@@ 2.6. Elements of potential theory
$\frac{{\rm d}\nu_{a}}{{\rm d}\theta}{(e^{{\rm i}\theta})}=\frac{1}{2\pi}\frac{\sin(\theta/2)}{\sqrt{\cos^{2}(\theta_{a}/2)-\cos^{2}(\theta/2)}},\quad e^{{\rm i}\theta}\in\Gamma_{a}.$ (2.36)
@@
$U_{\nu_{a}}(z)=-\int_{\theta_{a}}^{2\pi-\theta_{a}}\log\big|z-e^{{\rm i}\theta}\big|{\rm d}\nu_{a}\big(e^{{\rm i}\theta}\big)=-\log|G_{a}(z)|,\quad z\in{\mathbb{C}}\setminus{\partial\mathbb{D}},$ (2.37)
@@
$\frac{{\rm d}\nu_{\Delta}}{{\rm d}\theta}(e^{{\rm i}\theta})=\frac{1}{\pi p}\frac{\left|\Delta^{\prime}(e^{{\rm i}\theta})\right|}{\sqrt{4-\Delta^{2}(e^{{\rm i}\theta})}}.$ (2.38)
The prime designates differentiation of $\Delta(e^{{\rm i}\theta})$ with respect to $\theta$.
@@
$U_{\nu_{\Delta}}(z)=-\int_{\mathcal{B}_{\Delta}}\log\big|z-e^{{\rm i}\theta}\big|{\rm d}\nu_{\Delta}\big(e^{{\rm i}\theta}\big)=-\frac{1}{p}\log\left|G_{\Delta}(z)\right|,$ (2.39)
@@ 2.7. Zero counting measure and root asymptotics
Let $\mathcal{C}^{(0)}$ be a CMV matrix with either constant or periodic Verblunsky coefficients as in Section 2.4 or  2.5, respectively. Let $\Delta$, $\mathcal{B}_{\Delta}$, and $\nu_{\Delta}$ denote the discriminant, the associated finite-gap set, and the equilibrium measure, respectively.
Let $\mathcal{C}(\alpha_{0},\alpha_{1},\ldots)$ be another CMV matrix that is a compact perturbation of $\mathcal{C}^{(0)}$. In other words, $\alpha_{n}$ either converges to some $\alpha$ or is asymptotically $p$-periodic. Let $\Phi_{n}(z)$ and $\Phi^{(\beta)}_{n}(z)$ be the corresponding OPUC and POPUC, see (2.7) and (2.8). Denote by $z_{1,n},\dots,z_{n,n}$ the zeros of $\Phi_{n}$ counted with multiplicity, and by $z^{(\beta)}_{1,n},\dots,z^{(\beta)}_{n,n}$ the zeros of $\Phi_{n}^{(\beta)}$.
The zero-counting measures of $\Phi_{n}$ and $\Phi_{n}^{(\beta)}$ are defined as the probability measures
$\nu_{n}=\frac{1}{n}\sum_{j=1}^{n}\delta_{z_{j,n}}$ (2.42)
and
$\nu^{(\beta)}_{n}=\frac{1}{n}\sum_{j=1}^{n}\delta_{z^{(\beta)}_{j,n}}$. (2.43)
It is well known that for the case when $\mathcal{C}^{(0)}$ is not the free CMV matrix, both zero-counting measure converge to the corresponding equilibrium meausure, that is,
@@
For the case when $\mathcal{C}^{(0)}$ is the free CMV matrix, that is, when
$\lim_{n\to\infty}\alpha_{n}=0,$ (2.45)
we still have $\operatorname*{w-lim\ }_{n\to\infty}\nu_{n}^{(\beta)}=\nu_{0}$, the normalized Lebesgue measure on $\partial\mathbb{D}$}. However, things are more subtle for $\nu_{n}$. If one additionally assumes that
$\lim_{n\to\infty}|\alpha_{n}|^{1/n}=A\in[0,1],$ (2.46)

### v2

@@ 2.3. CMV matrices
From the spectral theory point of view, the probability measure $\mu$ associated with the sequence of Verblunsky coefficients $\{\alpha_{n}\}_{n=0}^{\infty}\subset{\mathbb{D}}^{\infty}$ is the spectral measure with respect to the cyclic vector $e_{0}$ of the unitary operator on $\ell^{2}(\mathbb{N}_{0})$, referred to as the CMV matrix:
$\mathcal{C}=\begin{pmatrix}\bar{\alpha}_{0}&\bar{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&0&0&\dots\\ \rho_{0}&-\bar{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&0&0&\dots\\ 0&\bar{\alpha}_{2}\rho_{1}&-\bar{\alpha}_{2}\alpha_{1}&\bar{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&\dots\\ 0&\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\bar{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&\dots\\ 0&0&0&\bar{\alpha}_{4}\rho_{3}&-\bar{\alpha}_{4}\alpha_{3}&\dots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}$, (2.5)
@@
Given a CMV matrix $\mathcal{C}$ associated with Verblunsky coefficients $\{\alpha_{n}\}_{n=0}^{\infty}$, the once-stripped CMV matrix, denoted by $\mathcal{C}^{(1)}$, is the CMV matrix constructed from the shifted sequence of Verblunsky coefficients $\{\alpha_{n}\}_{n=1}^{\infty}$; that is, the sequence obtained by removing $\alpha_{0}$ and reindexing. Similar definition applies to $\mathcal{C}^{(j)}$, which is the CMV matrix constructed from $\{\alpha_{n}\}_{n=j}^{\infty}$, for any $j\in\mathbb{N}_{0}$.
Let us also introduce the cut-off CMV matrix $\mathcal{C}_{n}$, which is the top-left $n\times n$ submatrix of matrix (2.5). It is easy to show that $\Phi_{n}$ is the characteristic polynomial of $\mathcal{C}_{n}$, i.e.
@@
An inspection of the structure of the CMV matrix shows that $\mathcal{C}_{n}$ depends only on the first $n$ Verblunsky coefficients. When needed, we will indicate this dependence by writing
@@
The POPUC (2.4) therefore satisfy
@@ 2.4. Carathéodory, Schur, and $m$-functions
$F(z)=\int_{0}^{2\pi}\frac{{\rm e}^{{\rm i}\theta}+z}{{\rm e}^{{\rm i}\theta}-z}{\rm d}{\mu({\rm e}^{{\rm i}\theta})}=\langle e_{0},(\mathcal{C}+z)(\mathcal{C}-z)^{-1}e_{0}\rangle,\quad z\in{\mathbb{D}}.$ (2.9)
@@
The Schur function is analytic $\mathbb{D}$} and satisfies $\sup_{z\in{\mathbb{D}}}|f(z)|\leq 1$. It will be convenient for us to extend (2.11) (and so also (2.15) below) to $z\in{\mathbb{C}}\setminus\mathop{\mathrm{supp}}\nolimits\,\mu$, which requires allowing $f(z)$ to take value $\infty$ at those points $z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}$ where F(z)=1F(z)=-1. For instance, under this convention, for the free CMV matrix ($\alpha_{n}\equiv 0$) we obtain $f(z)=0$ for all $z\in{\mathbb{D}}$ and $f(z)=\infty$ for all $z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}$, see Section 2.5. From (2.10), we obtain the symmetry
@@
Finally, we introduce a notation for the $(0,0)$-entry of the resolvent of $\mathcal{C}$:
$m(z)=\int_{0}^{2\pi}\frac{{{\rm d}{\mu({\rm e}^{{\rm i}\theta})}}}{z-{\rm e}^{{\rm i}\theta}}=\langle e_{0},(z-\mathcal{C})^{-1}e_{0}\rangle,\quad z\in{\mathbb{C}}\setminus\mathop{\mathrm{supp}}\nolimits\,\mu.$ (2.13)
@@
Formulas (2.14)–(2.16) are immediate from the definitions. Next, since $f$ does not take infinite values in $\overline{{\mathbb{D}}}\setminus\mathop{\mathrm{supp}}\nolimits\mu$, it follows from (2.12) that $f(z)\neq 0$ for $z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}$ as well as for $z\in{\partial\mathbb{D}}\setminus\mathop{\mathrm{supp}}\nolimits\mu$. Combining this with the first equality in (2.15) implies claim (ii). ∎
@@ 2.5. Constant Verblunsky coefficients
In the so-called free case, when $\alpha=0$, we have $\Phi_{n}(z)=z^{n}$ and the corresponding measure of orthogonality is the Lebesgue measure on $\partial\mathbb{D}$}. For $\alpha\in{\mathbb{D}}\setminus\{0\}$, the orthogonal polynomials $\Phi_{n}$ are known as the Geronimus polynomials, see [geronimus_vkgu66], as well as [golinski_maa99, Sec. 2] or [simon_opuc1, Ex. 1.6.12].
@@
where $\mathop{\mathrm{diag}}\nolimits$ denotes a diagonal matrix. The spectral measure $\mu$ of $\mathcal{C}(\alpha)$ with respect to $e_{0}$ is the Dirac delta measure $\delta_{\bar{\alpha}}$. This measure is trivial, and therefore it does not uniquely determine its orthogonal polynomials. Nevertheless, all the functions discussed below are unambiguously defined and consistent with this setting $\alpha\in{\partial\mathbb{D}}$, $\mu=\delta_{\bar{\alpha}}$.
Next we denote $a=|\alpha|\in[0,1]$,
@@
$\Gamma_{a}=\left\{{\rm e}^{{\rm i}\theta}\in{\partial\mathbb{D}}\;\big|\;\theta_{a}\leq\theta\leq 2\pi-\theta_{a}\right\}.$ (2.18)
@@
$\sigma_{\alpha}=\begin{cases}\Gamma_{a},&\mbox{ if }|\alpha+\tfrac{1}{2}|\leq\tfrac{1}{2},\\[2.0pt] \Gamma_{a}\cup\{\frac{1+\bar{\alpha}}{1+\alpha}\},&\mbox{ if }|\alpha+\tfrac{1}{2}|>\tfrac{1}{2}.\end{cases}$
@@
$\sqrt{\left(z-{\rm e}^{{\rm i}\theta_{a}}\right)\left(z-{\rm e}^{-{\rm i}\theta_{a}}\right)}=\sqrt{(z-1)^{2}+4za^{2}},$ (2.19)
@@
$\lim_{z\to 0}\sqrt{\left(z-{\rm e}^{{\rm i}\theta_{a}}\right)\left(z-{\rm e}^{-{\rm i}\theta_{a}}\right)}=1.$
@@
$G_{a}(z)=\frac{1}{2}\left(z+1+\sqrt{\left(z-{\rm e}^{{\rm i}\theta_{a}}\right)\left(z-{\rm e}^{-{\rm i}\theta_{a}}\right)}\right).$ (2.20)
@@
    $G_{0}(z)=\begin{cases}1,&\mbox{ if }z\in{\mathbb{D}},\\ z,&\mbox{ if }z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}},\end{cases}\qquad$ $F_{0}(z)=\begin{cases}1,&\mbox{ if }z\in{\mathbb{D}},\\ -1,&\mbox{ if }z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}},\end{cases}$
    $m_{0}(z)=\begin{cases}0,&\mbox{ if }z\in{\mathbb{D}},\\ \frac{1}{z},&\mbox{ if }z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}},\end{cases}\qquad$ $f_{0}(z)=\begin{cases}0,&\mbox{ if }z\in{\mathbb{D}},\\ \infty,&\mbox{ if }z\in{\mathbb{C}}\setminus\overline{{\mathbb{D}}}.\end{cases}$
@@
We prove part (iii). The expression (2.23) for $F_{\alpha}$ can be found in [golinski_maa99, Eq. (2.5)] (note that our $\alpha$ and $G_{a}$ correspond to their $\bar{a}$ and $z_{1}$, respectively). Combining (2.23) with (2.13) implies (2.24). Finally, expression (2.25) appears in [simon_opuc1, Eq. (1.6.82)]. Alternatively, one can deduce (2.25) by employing (2.11) and (2.22). ∎
@@ 2.6. Periodic OPUC
$\mathcal{B}_{\Delta}=\Delta^{-1}([-2,2])$
@@
In particular, the case of constant Verblunsky coefficients $\alpha_{n}\equiv\alpha\in{\mathbb{D}}$ from the previous section corresponds to
@@ 2.7. Elements of potential theory
$\frac{{\rm d}\nu_{a}}{{\rm d}\theta}{({\rm e}^{{\rm i}\theta})}=\frac{1}{2\pi}\frac{\sin(\theta/2)}{\sqrt{\cos^{2}(\theta_{a}/2)-\cos^{2}(\theta/2)}},\quad{\rm e}^{{\rm i}\theta}\in\Gamma_{a}.$ (2.28)
@@
$U_{\nu_{a}}(z)=-\int_{\theta_{a}}^{2\pi-\theta_{a}}\log\big|z-{\rm e}^{{\rm i}\theta}\big|{\rm d}\nu_{a}\big({\rm e}^{{\rm i}\theta}\big)=-\log|G_{a}(z)|,\quad z\in{\mathbb{C}}\setminus{\partial\mathbb{D}},$ (2.29)
@@
$\frac{{\rm d}\nu_{\Delta}}{{\rm d}\theta}({\rm e}^{{\rm i}\theta})=\frac{1}{\pi p}\frac{\left|\Delta^{\prime}({\rm e}^{{\rm i}\theta})\right|}{\sqrt{4-\Delta^{2}({\rm e}^{{\rm i}\theta})}}.$ (2.30)
The prime designates differentiation of $\Delta({\rm e}^{{\rm i}\theta})$ with respect to $\theta$.
@@
$U_{\nu_{\Delta}}(z)=-\int_{\mathcal{B}_{\Delta}}\log\big|z-{\rm e}^{{\rm i}\theta}\big|{\rm d}\nu_{\Delta}\big({\rm e}^{{\rm i}\theta}\big)=-\frac{1}{p}\log\left|G_{\Delta}(z)\right|,$ (2.31)
@@ 2.8. Zero counting measure and root asymptotics
Let $\mathcal{C}^{(0)}$ be a CMV matrix with either constant or periodic Verblunsky coefficients as in Section 2.5 or  2.6, respectively. Let $\Delta$, $\mathcal{B}_{\Delta}$, and $\nu_{\Delta}$ denote the discriminant, the associated finite-gap set, and the equilibrium measure, respectively.
Let $\mathcal{C}(\alpha_{0},\alpha_{1},\ldots)$ be another CMV matrix that is a compact perturbation of $\mathcal{C}^{(0)}$. In other words, $\alpha_{n}$ either converges to some $\alpha$ or is asymptotically $p$-periodic. Let $\Phi_{n}(z)$ and $\Phi^{(\beta)}_{n}(z)$ be the corresponding OPUC and POPUC, see (2.6) and (2.7). Denote by $z_{1,n},\dots,z_{n,n}$ and $z^{(\beta)}_{1,n},\dots,z^{(\beta)}_{n,n}$ the zeros of $\Phi_{n}$ and $\Phi_{n}^{(\beta)}$, respectively. Recall the definition of the zero-counting measures of $\Phi_{n}$ and $\Phi_{n}^{(\beta)}$,
[…]
It is well known that for the case when $\mathcal{C}^{(0)}$ is not the free CMV matrix, both zero-counting measures converge to the corresponding equilibrium meausure, that is,
@@
For the case when $\mathcal{C}^{(0)}$ is the free CMV matrix, i.e., $\alpha_{n}\to 0$, we still have $\nu_{n}^{(\beta)}\to\nu_{0}$ weakly as $n\to\infty$, where $\nu_{0}$ the uniform probability measure on $\partial\mathbb{D}$}. However, things are more subtle for $\nu_{n}$. If one additionally assumes that
$\lim_{n\to\infty}|\alpha_{n}|^{1/n}=A\in[0,1],$